OpenStudy (anonymous):

suppose a parabola has vertex (5, –3) and also passes through the point (6, 1). Write the equation of the parabola in vertex form. y = (x – 5)2 – 3 y = 4(x – 5)2 – 3 y = 4(x + 5)2 – 3 y = 4(x – 5)2 + 3

OpenStudy (tkhunny):

Since y = x^2 is the only form offered, we can ignore the x = y^2 variety. Should be an immediate result this far: (y-(-3)) = (something)(x-5)^2 This discards #3. It has "x+5" Standard form (y+3) = (something)(x-5)^2 y = (something)(x-5)^2 - 3 This discards #4. It has "+3". Substitute (6,1) 1 = (something)(6-5)^2 - 3 1 = (something)(1)^2 - 3 1 = (something) - 3 Something = 4 Discard #1 and we're done. Really, just one step at a time. Walk through it in a way that is logical and organized. No need to see the solution from the beginning, Just take a step.

OpenStudy (anonymous):

so its y = 4(x – 5)2 – 3

OpenStudy (tkhunny):

Please use the caret for exponentiation. x2 is very odd-looking. x^2 means x-squared. You tell me. Is it?